OpenStudy (anonymous):

log (x+3) = 1 - log x

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

\[\log(x+3)+\log(x)=1\] \[\log(x(x+3))=1\] \[x(x+3)=10\] \[x^2+3x-10=0\] etc

OpenStudy (anonymous):

the "etc" is \[(x+5)(x-2)=0\] so \[x=2\] because x cannot be 5 because you cannot take the log of a negative number